OpenStudy (anonymous):

is this correct ? what are the solutions ? 16x^2+9=0 A. -3/4i,3/4i B. -4/3i,4/3i C. -3/4, 3/4 <----- my answer D. -9/16i,9/16i

OpenStudy (anonymous):

if you plug those back in for x do you get the original problem back?

OpenStudy (anonymous):

what do you mean ?

OpenStudy (anonymous):

maybe it would be better if I asked you how you got that answer

OpenStudy (anonymous):

16x^2-9=0 (4x+3)(4x-3)=0 4x+3=0; 4x-3=0 4x=-3; 4x=3 x=-3/4; x=3/4

OpenStudy (anonymous):

try subtracting 9 from each side and then taking the square root of each side. What do you get?

OpenStudy (anonymous):

your original problem says 16x^2+9 not -9, which is correct?

OpenStudy (anonymous):

oh lol my bad